OpenStudy (cloverracer):

PLEASE HELP WAITED 1 HR FOR HELP! :( PB has endpoints at -2 and 12. What is the coordinate of its midpoint? (Whoever explains, can you please show steps so I can fully understand? Thanks!)

OpenStudy (acxbox22):

wait.....a point does not have a midpoint

OpenStudy (gschibby):

Again, what is PB?

OpenStudy (cloverracer):

It is what my homework says :/

OpenStudy (gschibby):

What's your subject/theme?

OpenStudy (cloverracer):

Geometry

OpenStudy (acxbox22):

a line has to have two endpoint coordinates

OpenStudy (cloverracer):

This is exactly what it says. PB has endpoints at -2 and 12. What is the coordinate of its midpoint?

OpenStudy (gschibby):

Exactly, what acxbox says, so I guess this line just goes along the x-axis ;) A line that goes from -2 to 12 is 12+2=14 in length, divide that by 2 and you have the center point ;)

OpenStudy (acxbox22):

no like two coordinates like (2,0) and (5,4)

OpenStudy (cloverracer):

\[12 + 2 = 14 / 2\] like this?

OpenStudy (cloverracer):

I have 7 then

OpenStudy (cloverracer):

@acxbox22 @Gschibby

OpenStudy (phi):

you found the distance to the mid point (it is 7 more than -2 and also 7 less than 12) you can find the actual mid point by doing -2 + 7 you could also find the mid point by find the average of the two end points: (-2 + 12)/2

OpenStudy (cloverracer):

So for the actual mid point I got 5

OpenStudy (phi):

yes. Here is a picture

OpenStudy (cloverracer):

Okay I have a solution am I correct?\[M = -2 + 12/2 = 5/2 = 2.5\]

OpenStudy (cloverracer):

The coordinate of the midpoint of PB is 2.5?

OpenStudy (phi):

M= (-2+12)/2 = 10/2 = 5 (not 2.5)

OpenStudy (gschibby):

Sorry phi, but I think you are wrong here. 7 would be the right answer if the line goes along the x-axis :)

OpenStudy (gschibby):

No, pellete, you're absolutely right! My bad! :P

OpenStudy (cloverracer):

so \[M = -2 + 12 / 2 = 10/2 = 5\] The coordinate of the midpoint of PB is 5

OpenStudy (phi):

yes

OpenStudy (gschibby):

Oops, bad language warning...sorry about that! :P

OpenStudy (cloverracer):

Thank you for your help! Can you help me on 2 more?